# The Neuromorphic Analog Electronic Nose

### v2

@@ 1 Introduction
For most pattern recognition methods, the sensor data has to be sampled so that it can be fed into a computer or microprocessor. ADCs typically employed in an e-nose system usually have a fixed sampling frequency. Therefore, the samples are generated at a fixed time interval throughout the operation of the e-nose system, irrespective of the presence or absence of any gas in the environment. This generates a large number of irrelevant samples to deal with at the pattern recognition stage which increases the computational cost. If we look at biological neural systems, the data is transmitted from one point to another only when there is some relevant information. This leads to significant data reduction and makes the inference easier.
@@ 2.1 Experimental Setup for E-nose Recordings
A custom-made e-nose [39] was used for data collection and it comprises four different MOx gas sensors operated at a constant heater voltage. Out of these four available sensors, reducing (RED) (Sensor 1), oxidizing (OX) (Sensor 2), and Ammonia ($NH_{3}$) (Sensor 3) sensors of MiCS-6814 have been used for our circuit design. The odor stimuli were provided by a multi-channel odor delivery device described in detail by Ackels et al. [40]. It offers an exceptionally high temporal fidelity, which is achieved by combining high-speed gas valves, flow controllers, as well as short and narrow gas pathways. Odorant headspace samples were embedded in a constant airflow and presented to the e-nose. The delivered odor concentration was varied by modulating the valve shatter duty cycle, ensuring a linear relationship between different concentration levels. Fluctuations in the flow during gas presentations were minimized by careful calibration. For our design, recordings of three different odor stimuli have been used: Ethyl Butyrate (EB), Eucalyptol (Eu), and Isoamyl Acetate (IA). These gases were diluted in an odorless mineral oil at a ratio of 1:5.
@@
Figure 1 shows the mean sensor measurements over 20 trials for three odor stimuli at five different concentration levels for three sensors of MiCS-6814. The measurement consists of voltage recordings across the load resistors connected in series with sensors. This voltage varied inversely with respect to MOx sensor resistance and was used as the input signal to the circuit. The load resistance values connected to sensors 1, 2 and 3 are $27k\Omega$, $6.8k\Omega$, and $27k\Omega$ respectively. Negative timestamps indicate the sensor’s baseline response before gas release.
@@ 3.1.1 The Change Detection Circuit
The change detection circuit consists of an inverting op-amp differentiator U1 followed by an op-amp comparator U2. The MOx sensor is connected in series with a load resistor. The voltage drop V1 across this load serves as an input to the circuit. The inverting op-amp differentiator U1 is provided with the DC offset at its inverting terminal which is adjustable by the potentiometer X1. The gain of the differentiator depends on resistor R2 and capacitor C1. All resistors and capacitors used in the circuit together determine the rise and decay time of the differentiator output.
The inverting differentiator output is compared with a pre-defined threshold voltage by the comparator U2. A change detection pulse ($Out_{CD}$) is generated whenever the output of the differentiator exceeds this threshold. This threshold voltage level is adjustable by the potentiometer X2 and can be varied depending on the SNR of the input signal.
@@
Figure 4 shows the output of the CD circuit for EB gas at concentration level 1. The red curve shows the inverting differentiator output. The time constants of the differentiator are chosen in such a way that it has a high gain during the rising flank of the input signal and it reverts as soon as the input signal starts falling. Therefore, the CD pulse $Out_{CD}$ remains high during the rising flank and resets as soon as the input signal switches sign.
@@ 3.2.1 The Timer Circuit
Figure 7 shows the output of the timer circuit for EB gas at concentration level 1. The black pulse is the $CD_{out}$ pulse which turns on the SR latch. Once the SR latch turns on, the latch output $Q_{out}$ goes high which charges the $C_{ramp}$ capacitor. The voltage across $C_{ramp}$ (shown by the blue curve) rises with capacitor charging. Once this voltage reaches the threshold level (which is 0.9 V here), the latch is reset by the comparator $U_{ramp}$, and the $Q_{out}$ pulse goes low. The start timing of the $Q_{out}$ pulse depends on the CD pulse generated by the fastest sensor. But its width is only decided by the charging time of the capacitor $C_{ramp}$ which is fixed in a particular setting and is independent of the CD pulses generated by individual sensors in the array. Therefore, the exposure measurement of all sensors triggered by the $Q_{out}$ pulse will start at the same time and its duration will be the same for all sensors irrespective of whether any sensor is fast or slow.
@@ 4.1 Single Sensor
The proposed circuit for the single sensor was simulated and differences in timings of activation of CD and EM pulses were measured for all gases at all five different concentration levels. Figure 8 shows the plot of the inverse of the mean time difference over all trials for CD and EM pulse activation with respect to the concentration level of each gas. Error bars at each concentration level represent the standard deviation over all 20 trials. It can be observed that the time difference between CD and EM pulses varies inversely with gas concentration. In this way, the gas concentration information is encoded in the timings of analog spikes generated from two separate pathways, analogous to the spiking output of the mammalian olfactory bulb.
@@ 5 Conclusion and Discussion
This work presents simple analog circuit implementations for encoding gas concentrations in the timings of spikes obtained on two parallel pathways, inspired by the mechanism of a mammalian olfactory bulb. These circuits generate output only when there is some significant change in the environment due to gas injection, which is in contrast to ADCs that sample the gas sensor data even when there is no useful information. Therefore, the proposed circuits for a single sensor and a sensor array can be efficient substitutes for ADCs that can be interfaced with a MOx sensor and sensor array respectively.
@@
Apart from the olfactory bulb, parallel processing streams also exist in the preprocessing stations of other sensory modalities such as the retina of the eye [48, 49, 50, 51], cochlea of the ear [52, 53, 54], and also in the mammalian tongue [55] and the skin [56, 57, 58]. These findings suggest that separating different stimulus features into separate channels at the receptor level before transmission into higher brain regions for processing seems to be a common and important property in all sensory modalities. Network simulations performed by Geramita et al. [59] prove that tufted cells perform best in odor discrimination at low concentration ranges and mitral cells perform best at high concentration ranges. The parallel combination of MCs and TCs discriminates odor best at intermediate concentration ranges and second best at low and high concentration ranges. This finding indicates that the parallel pathway system encodes some new information that is unlikely to occur in single pathway systems.
@@
The analog circuits for a single gas sensor and the gas sensor array have been tested for MOx recordings which have been collected over 12 hours. Therefore, the trial-to-trial variability observed in these recordings due to the interaction of chemical analytes at the sensing film because of long-term usage of sensors (first-order sensor drift [60]) is small. Also, these recordings have been obtained in a controlled environment setting with a regular airflow and therefore have not been impacted by the changes in environmental conditions, or noise in the measurement delivery system (second-order drift [60]).
@@
Long-term exposure to gases often makes MOx sensors prone to first-order drift, which shifts the baseline resistance of MOx sensors [61]. Considering long-term experiments, we expect that our proposed circuits are resilient to this first-order baseline drift because the circuits use the differentiator as one of the subcircuits to cancel out this baseline shift. Additional long-term experiments are needed to fully assess how drift affects these circuits. Our conviction is that the MOx array circuit is much more robust in terms of drift effects than the single sensor circuit. This is because the circuit readouts from other cleaner sensors in the array would compensate for the drift-ridden sensor.
@@
The proposed circuits can be used in a scenario where sensors are exposed to a single gas pulse of varying concentrations. However, a more realistic environment contains gas plumes that are formed as gas molecules are dispersed by the wind away from the source [62]. In such an environment, the gas concentration is not fixed but changes dynamically in both spatial and temporal domains because of turbulence [63]. Therefore, the response of a MOx sensor in these environments could be very different from that in a controlled single pulse environment [63, 64, 65]. It would be of interest in the future to analyze the MOx sensor data recorded for more natural gas plumes with time-varying concentrations and to do further design modifications in the proposed circuits if needed to make them work properly in more realistic environments.

### v3

# A Neuromorphic Electronic Nose Design

@@ 1 Introduction
For most pattern recognition methods, the sensor data has to be sampled to be used by a computer or microprocessor. ADCs typically employed in an e-nose system usually have a fixed sampling frequency. Therefore, the samples are generated at a fixed time interval throughout the operation of the e-nose system, irrespective of the presence or absence of any gas in the environment. This generates a large number of irrelevant samples to deal with at the pattern recognition stage which increases the computational cost. In biological neural systems, data is transmitted between two nodes only when it carries relevant information. This leads to significant data reduction and makes the inference easier.
@@ 2.1 Experimental Setup for E-nose Recordings
The data collection campaign using a custom-made e-nose has been described in a previous study [39], and leverages four different MOx gas sensors operated at a constant heater voltage. Out of these sensors, reducing (RED) (Sensor 1), oxidizing (OX) (Sensor 2), and Ammonia ($NH_{3}$) (Sensor 3) sensors of MiCS-6814 have been used for our circuit design. For the data collection, the odor stimuli were provided by a multi-channel odor delivery device described in detail by Ackels et al. [40]. It offers an exceptionally high temporal fidelity, which is achieved by combining high-speed gas valves, flow controllers, as well as short and narrow gas pathways. Odorant headspace samples were embedded in a constant airflow and presented to the e-nose. The delivered odor concentration was varied by modulating the valve shatter duty cycle, ensuring a linear relationship between different concentration levels. Fluctuations in the flow during gas presentations were minimized by careful calibration. For our design, recordings of three different odor stimuli have been used: Ethyl Butyrate (EB), Eucalyptol (Eu), and Isoamyl Acetate (IA). These gases were diluted in an odorless mineral oil at a ratio of 1:5.
@@
Figure 1 shows the mean sensor measurements over 20 trials for three odor stimuli at five different concentration levels for three sensors of MiCS-6814. The relative gas concentration at these concentration levels was 20 % (C1), 40 % (C2), 60 % (C3), 80 % (C4), and 100 % (C5). The measurement consists of voltage recordings across the load resistors connected in series with sensors. This voltage varied inversely with respect to MOx sensor resistance and was used as the input signal to the circuit. The load resistance values connected to sensors 1, 2 and 3 are $27k\Omega$, $6.8k\Omega$, and $27k\Omega$ respectively. Negative timestamps indicate the sensor’s baseline response before gas release.
@@ 3.1.1 The Change Detection Circuit
The change detection circuit consists of an inverting op-amp difference circuit U1 with a band-pass filtering operation followed by an op-amp comparator U2. The MOx sensor is connected in series with a load resistor. The voltage drop V1 across this load serves as an input to the circuit. The inverting op-amp difference circuit U1 is provided with the DC offset at its non-inverting terminal which is adjustable by the potentiometer X1. The gain of the circuit depends on resistor R2 and capacitor C1. All resistors and capacitors used in the circuit together determine the rise and decay time of the circuit output.
The inverting difference circuit output is compared with a pre-defined threshold voltage by the comparator U2. A change detection pulse ($Out_{CD}$) is generated whenever the output of the differentiator exceeds this threshold. This threshold voltage level is adjustable by the potentiometer X2 and can be varied depending on the SNR of the input signal.
@@
Figure 4 shows the output of the CD circuit for EB gas at concentration level 1. The red curve shows the inverting filtered difference circuit output. The time constants of the circuit are chosen in such a way that it has a high gain during the rising flank of the input signal and it reverts as soon as the input signal starts falling. Therefore, the CD pulse $Out_{CD}$ remains high during the rising flank and resets as soon as the input signal switches sign.
@@ 3.2.1 The Timer Circuit
Figure 7 shows the output of the timer circuit for EB gas at concentration level 1. The black pulse is the $CD_{out}$ pulse which turns on the SR latch. Once the SR latch turns on, the latch output $Q_{out}$ goes high which charges the $C_{ramp}$ capacitor. The voltage across $C_{ramp}$ (shown by the blue curve) rises with capacitor charging. Once this voltage reaches the threshold level (which is 0.9 V here), the latch is reset by the comparator $U_{ramp}$, and the $Q_{out}$ pulse goes low. The start timing of the $Q_{out}$ pulse depends on the CD pulse generated by the fastest sensor. But its width is only decided by the charging time of the capacitor $C_{ramp}$ which is fixed in a particular setting and is independent of the CD pulses generated by individual sensors in the array. Therefore, the exposure measurement of all sensors triggered by the $Q_{out}$ pulse will start at the same time and its duration will be the same for all sensors irrespective of whether any sensor is fast or slow. As the start timing of the exposure measurement pulse will be different for different sensors depending on their sensitivities to a given gas, the gas concentration information from each sensor is encoded in the timing of the exposure measurement pulse relative to the timing of the global trigger pulse.
@@ 4.1 Single Sensor
The proposed circuit for the single sensor was simulated and differences in timings of activation of CD and EM pulses were measured for all gases at all five different concentration levels. Figure 8 shows the plot of the inverse of the mean time difference over all trials for CD and EM pulse activation with respect to the concentration level of each gas. Error bars at each concentration level represent the standard deviation over all 20 trials. It can be observed that the time difference between the rising edges of CD and EM pulses varies inversely with gas concentration. In this way, the gas concentration information is encoded in the timings of analog spikes generated from two separate pathways, analogous to the spiking output of the mammalian olfactory bulb. To convert the finite duration analog pulses into short spikes as observed in biology, we can pass the CD and EM pulses through an edge detector circuit that generates fixed-duration spikes only at the positive edge of these pulses. These spikes can then be transmitted at the inference stage.
@@ 5 Conclusion and Discussion
This work presents simple analog circuit implementations for encoding gas concentrations in the timings of spikes obtained on two parallel pathways, inspired by the mechanism of a mammalian olfactory bulb. These circuits generate output only when there is some significant change in the environment due to gas injection, which is in contrast to ADCs that sample the gas sensor data even when there is no useful information. Therefore, the proposed circuits for a single sensor and a sensor array can be data-efficient substitutes for ADCs that can be interfaced with a MOx sensor and sensor array respectively.
@@
Apart from the olfactory bulb, parallel processing streams also exist in the preprocessing stations of other sensory modalities such as the retina of the eye [48, 49, 50, 51], cochlea of the ear [52, 53, 54], and also in the mammalian tongue [55] and the skin [56, 57, 58]. These findings suggest that separating different stimulus features into separate channels at the receptor level before transmission into higher brain regions for processing seems to be a common and important property in all sensory modalities. Network simulations performed by Geramita et al. [59] prove that tufted cells perform best in odor discrimination at low concentration ranges and mitral cells perform best at high concentration ranges. The parallel combination of MCs and TCs discriminates odor best at intermediate concentration ranges and second best at low and high concentration ranges. This finding indicates that the parallel pathway system facilitates odor discrimination across a wider range of concentrations as compared to single pathway systems.
The circuits described in this work currently require manual adjustment of potentiometers for comparator thresholds, which could be cumbersome for a large sensor array. It would be of interest in the future to investigate some automated calibration strategies [60] [61] for adaptive comparator thresholds, which will reduce the manual operation and hence the operation of these circuits can be scaled for bigger e-nose systems.
@@
The analog circuits for a single gas sensor and the gas sensor array have been tested for MOx recordings, which have been collected over 12 hours. Therefore, the trial-to-trial variability observed in these recordings due to the interaction of chemical analytes at the sensing film, because of long-term usage of sensors (first-order sensor drift [62]) is small. Also, these recordings have been obtained in a controlled environment setting with a regular airflow and therefore have not been impacted by the changes in environmental conditions, or noise in the measurement delivery system (second-order drift [62]).
@@
Long-term exposure to gases often makes MOx sensors prone to first-order drift, which shifts the baseline resistance of MOx sensors [63]. Considering long-term experiments, we expect that our proposed circuits are resilient to this first-order baseline drift because the circuits use the band-pass filter as one of the subcircuits to cancel out this baseline shift. Additional long-term experiments are needed in the future to fully assess how drift affects these circuits. Our conviction is that the MOx array circuit is much more robust in terms of drift effects than the single sensor circuit. This is because the circuit readouts from other cleaner sensors in the array would compensate for the drift-ridden sensor.
The proposed sensor array front-end has been designed and tested only for a 3-sensor array. To further validate the robustness of this circuit to drift effects and to deploy it in complex environments comprising other interfering gases apart from the target gas, the circuit should be tested for a large sensor array comprising MOx sensors of a diverse range of sensitivities to a particular gas. However, scaling up the circuit for large arrays would create additional issues such as less area efficiency due to timer circuit and OR gates, and false positives due to cross-sensitivity of gas sensors [64]. Reducing the circuit complexity for large array systems to create area-efficient e-nose front-ends and further circuit modifications for reducing the number of false positives are valuable future directions for this work.
@@
The proposed circuits are environment-sensitive and can only be used in a scenario where sensors are exposed to a single gas pulse of varying concentrations in constant airflow. However, a more realistic environment contains gas plumes that are formed as molecules of a single gas or a mixture of gases are dispersed by the wind away from the source [65]. In such an environment, the gas concentration is not fixed but changes dynamically in both spatial and temporal domains because of turbulence [66]. Therefore, the response of a MOx sensor in these environments could be very different from that in a controlled single pulse environment [66, 67, 68]. It would be of interest in the future to analyze the MOx sensor data recorded for more natural plumes of a single gas or gas mixtures with time-varying concentrations and to do further design modifications in the proposed circuits [69] if needed to make them work properly in more realistic environments.